Shift report 20000725 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 25 2000 - 16:16:20 EDT

  • Next message: http account: "Shift report 20000725 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Jul 25 16:16:20 2000, Shift: 08:00-16:00
    Supervisor: YKL, with: EJK
    Log book#3 pages 76 - 78
    
    Shift Summary:
    ------------------------------------------------------------
    The beams were cogged at 70 G till 11AM. The beam was
    steered to PHENIX at 10AM and collision rate was lowered
    for BRAHMS. While the beams were steered in BRAHMS IR
    as a response to our complaints the beams were lost.
    Quench interlock and PS problems occured and beams did
    not come back.
    
    Run 1929 was continued for a total of 4 and 1/2 hourse
    with trigger 6 counts of 2K. FV worked on ZDC trigger, and 
    changed TPM1 gating grid.
    
    Memory leaks seem to be related to "TPCPreprocess".
    Without "TPCPreprocess", SuperMon was running smoothly.
    FV Updated operator account for online monitor 
    with ZDC changes. 
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 25 2000 - 16:20:02 EDT